Shift Summary May 29 DAY shift

Vivian O'Dell (odell)
Sat, 29 May 99 16:01:08 -0500

Shift Summary for Day shift, 05/29/99
On Shift: Albert Erwin
Vivian O'Dell
A variety of other committed collaborators were working throughout this
shift.

Hours of beam delivered at all intensities: 0.1666667
Hours of beam delivered at nominal intensity: 0.1666667
Hours of 99 data (special runs not included): 0
Hours of special runs (laser scans, muon runs...): 0.0 hours

Accounting of lost time:

No real beam this shift due to TeV studies.

Activities:

7:15 Hogan and Theo gave up on replateauing DC's -- made an access
to lower threshold back to 40 mV.

7:30 Al & Viv made an access. Under superb supervision from
Rick T (who remained upstairs), Al replaced the terminator on
the PMT base for MU3 channel 32X which fixed that channel.

7:30 took loooooooooooong laser scan 13215.

7:30 Sydney finds a problem in the onsplit s/w and fixes it.
Another problem he couldn't fixed and sent email to JAG.

10:00 Al realized he forgot his sweater downstairs. He remained
cold the rest of the shift.

12:00 Al made more copies of the gas check list, and checks the
gas. He notes that the alcohol level can no longer be checked because
of the backpressure on the bubbler.

14:00 Arun tests new damp_summary. Mostly works.

14:20 Many CsI channels have shifted quite a bit, according to the
muon run from last night (see list in logbook). Ed installed new
CsI Q/E constants based on muon run 13208.

15:15 Albert and Valeri made an access to adjust HV on V/V'. Albert
retrieves his sweater. Interlocks accidently dropped. Call MCR. They
will send someone out for a search & secure but don't know when.

15:30 new shift takes over.

=========================================================================
Runs ended this shift:
Run 131215 - loooong laser scan

Tapes brought to FCC this shift: 0